the sides of a triangle are y, y+1 and 7 centimeters,if the perimeter is 56 centimeters, what is the value of y
Answer by jim_thompson5910(35256)

The perimeter is the sum of the sides. So

P = A+B+C
P = y+(y+1)+7
P = 2y + 8

But we're given the perimeter of 56. So

56 = 2y+8 or 2y + 8 = 56

Solve this equation for y to get your answer.